Two samples, one of size 14 and the second of size 13, are selected to test the difference between two population means. How many degrees of freedom are used to find the critical value? Assume the population standard deviations are equal.
Answer

27

26

25

14

Answer by stanbon(57967)

Two samples, one of size 14 and the second of size 13, are selected to test the difference between two population means. How many degrees of freedom are used to find the critical value? Assume the population standard deviations are equal.
Answer
27
26
25
14
-----------------
df = n1+n2-2 = 14+13-2 = 25
==================================
Cheers,
Stan H.
=====================